# FoodieQA: A Multimodal Dataset for Fine-Grained Understanding of Chinese Food Culture

### v1

@@ 1 Introduction
Figure 2: The tasks in FoodieQA evaluate food culture understanding from three perspectives. Multi-image VQA requires the ability to compare multiple images, similar to how humans browse a restaurant menu. Single-image VQA assesses whether models can use visual information to better understand food culture. Text-based questions probe model performance without multimodal data. Fine-grained attributes that the questions focus on are highlighted.
To build a regionally diverse dataset, we gather dishes and images selected by native Chinese speakers from various regions, covering 14 distinct cuisine types across China. To ensure the images used for benchmarking are fresh and have no chance of leaking into the pretraining data of VLMs, we collect images uploaded by local people, which are not publicly available online. We then define multiple attributes associated with the dishes and have native Chinese annotators create multiple-choice questions based on their expertise. Our dataset includes both text-based question answering and vision-based question answering tasks, as illustrated in Figure 2.
@@ 2 Related Work
In recent years, most food datasets have been designed for food image classification (Chen et al., 2017), food captioning (Ma et al., 2023), and recipe-focused generation and question answering (Yagcioglu et al., 2018; Min et al., 2018; Liu et al., 2022). For culture knowledge probing in the food domain, some of the recent datasets span multiple countries and include broad cultural or regional metadata (Min et al., 2018; Ma et al., 2023). However, they often use country as a proxy for culture, such as the country of origin for the food. For example, Palta and Rudinger (2023) introduced a test set to probe culinary cultural biases by considering US and non-US traditions, and Cao et al. (2024a) focuses on recipe transfer between Chinese and English. Investigating cultural differences within a country remains an under-explored area (Palta and Rudinger, 2023).
@@
Bugliarello et al. (2023) quantified the fine-grained vision-language understanding capabilities in existing models, focusing on aspects within the general domain. Later works focus on the culture understanding in VL models Liu et al. (2023); Cao et al. (2024b). However, current fine-grained VL datasets (Zhang et al., 2021; Parcalabescu et al., 2022; Thrush et al., 2022; Hendricks and Nematzadeh, 2021) are often framed as binary classification tasks, which limits their difficulty. Our multi-choice vision question answering dataset aims to advance the boundaries of fine-grained understanding in the context of food and culture.
@@ 3.3 Local Specialty Annotation
We also gather text annotations of representative local specialties for each cuisine type on our list. Annotators are asked to collect meta information for representative local dishes for each cuisine type, based on their life experience and knowledge obtained from the web. These meta-fields provide information beyond recipes, offering insights into how the food looks and tastes when people are eating it. An example is provided in Figure 4. The annotation is done by eight native Chinese speaker which include five PhD students and three postdoctoral researchers from different provinces in China.
The 17 meta-info fields cover the looks, taste, and culinary attributes of a dish. They include the food category, dish name, alternative names, main ingredient, characteristics of the main ingredient, three other key ingredients, dish flavor, presentation style, dish color, serving temperature (cold or warm), dishware used, region and province of origin, cuisine type, three primary cooking techniques, eating habits (if any), and reference links.
@@ 3.4.1 Multi-image VQA
Once the questions and answers for the multi-image multiple-choice questions are collected, we verify the questions by asking the annotators (who did not create the questions) to answer them. If a question does not meet our defined criteria, annotators are instructed to flag it as a "bad question." Through this process, 87 questions were discarded. Additionally, when answering the questions, annotators are required to provide the rationale they use to reach the answer, as well as judge whether the question requires multi-hop reasoning. The user interface that we use for verification is shown in Figure 13. Each question is verified by two annotators, and we exclude the questions that do not have full agreement.
@@ 3.4.2 Single-Image VQA
Besides using images as multiple-choice answer options, we also ask diverse fine-grained questions about various aspects of a dish based on its meta-information. We identify dishes that have both meta-information annotations and collected images, and then create questions based on the meta-information. As shown in the example in Figure 2, the dish name is intentionally omitted from the questions to ensure they can only be answered by examining the visual features.
@@
We adopt a template-based approach, where a question about the same meta-field is asked multiple times, varying factors like the image of the dish, while the answer options are carefully selected from the wrong candidates in the meta-field to ensure that only one answer is correct. The single-image VQA questions are generated using a rule-based method, followed by thorough human verification and filtering through that is similar to the multi-image VQA verification process. Please see details in the Appendix A.
@@
Similar to verification for the multi-image VQA questions, annotators are asked to answer the question given the text question and the corresponding image, and raise a “bad-question” flag to filter out questions that does not satisfy the criteria. 88 questions were discarded as bad. Note that the name of the dish is not revealed in the text question so that the question needs to be answered based on visual information. Annotators are asked to write “I don’t know" in the rationale and randomly guess an answer if they think the question is beyond their knowledge.
@@ 3.5 Text Question Answering Annotation
We formulate the text-based questions by combining human annotations and rule-based generation. Similar to the single-image VQA approach described in Section 3.4.2, we generate questions and multiple-choice answer options based on the meta-information fields. However, instead of using the dish image, we included the dish name directly in the question. The questions are formulated using templates, where only the dish names and meta-fields are varied. A same human verification process to single-image question answering is included. 135 bad questions are discarded. Notice that annotators are asked to answer the questions based on their knowledge without using search engines, this makes the task challenging as it would be hard for one to answer questions about unfamiliar foods and regions without any other available information besides names of the food.
@@ 4.1 Human Validation
In Table 1, we calculate human accuracy and inter-annotator agreement scores based on human-verified questions, excluding those identified as bad questions. For the single-image VQA and text QA questions, given the diverse cultural backgrounds of the human annotators, some questions can be challenging if the required food culture knowledge falls outside an annotator’s cultural experience. Annotators are instructed to indicate "I don’t know" when they lack the cultural knowledge to answer a question. These questions are classified as out-of-domain. For out-of-domain questions, the answer is randomly selected from the provided choices when calculating human accuracy and Cohen’s Kappa scores. We also report Cohen’s Kappa ($\kappa$), and human accuracy for in-domain questions. The human validation process involves eight native Chinese speakers from seven different provinces across China [^4] The annotators are from Sichuan, Shaanxi, Guangdong, Jiangsu, Jiangxi, Shandong, and Chongqing., including three postdoctoral researchers and five PhD students. Each question is verified and answered by two annotators.
Figure 5: Region distribution of collected food images.
@@
Table 1: Question Statistics per task in FoodieQA.
@@ 4.2 Image and Question Distribution
After human verification, we obtain 403 multi-image VQA questions, where each question needs to be answered with a set of four provided images. Single-image VQA tasks consists of 256 question in total, and text QA consists of 705 questions in total (Table 1). We report the key statistics of the questions in Table 2. Please see more details in Appendix B.
@@ 5 Baselines: How Much of a Foodie are the LLMs/VLMs?
We evaluate open-weight and API-based state-of-the-art LLMs and VLMs to probe their culture knowledge in the food domain. We evaluate the models in both Chinese and English [^5] We also include an estimate, calculated over 100 random samples, of Human performance on the English Multi-Image and Single-Image VQA from one native speaker with no specialized knowledge of Chinese food culture.. For VQA questions are translated to English using the DeepL free API [^6] https://www.deepl.com/en/translator and validated by two PhD students.
@@ 5.1 Multi-Image VQA is Difficult
We experimented with four different prompts that utilized lists of images and texts or interleaved image-text inputs. Details can be found in Appendix C. As shown in Figure 6, when compared to the human accuracy of 91.69% in Chinese, the best-performing open-weight model, Idefics2-8B, achieves an accuracy of 50.87%, which is still significantly lower than human performance. This indicates that current state-of-the-art models are still weak at distinguishing differences among food from visual input. This underscores that multi-image understanding, especially in contexts requiring cultural knowledge in the food domain, remains a challenging problem. When evaluating on the translated English questions, model performance decreases for all models except Phi-3-vision.
@@ 5.2 Single-Image VQA Results
Table 3: Comparison of Multi-image and Single-image VQA Performance in Chinese and English. We report the best accuracy from four prompts. : see Footnote 4.
@@ 5.3 Models are Strong at Text QA
From Figure 7, we see that the Qwen2-7B-instruct model surpasses human performance on the text QA task, where the questions are formulated based on the local specialty annotations in Section 3. Since the local specialty annotations are collected and summarized from public resources such as Baidu-Baike by local representatives, we suspect that the high performance could be due to the inclusion of domain-specific training data.
@@ 6 Analysis
(a) Accuracy by food categories.
(b) Accuracy by cuisine types.
@@
In Single-image VQA, the default setting is to use only dish image without specifying the dish name. We now examine whether the visual information is beneficial using the Idefics2-8B model. [^7] We selected this model because it supports text-only inputs, unlike some other models such as the Yi-VL series. Results are shown in Table 4, where we investigate two variants of Single-image VQA: providing the model with dish name only versus both the dish name and image. We observe that the Idefics2 model consistently performs better when dish images are available as additional information. Please see comparison examples in Appendix E.2.
@@ 7 Conclusion
Improved models or methods for understanding food culture may be essential for future progress in the FoodieQA challenge. Looking ahead, we aim to expand the dataset to include dishes from other countries and regions. We make all of our data collection, annotation, and verification tools freely available for re-use, and encourage the community to create Foodie datasets for their own language. [^8] We will release our dataset as a benchmark on Codabench.
@@ Appendix A Rule-based question formulation
For text-based question answering we develop a rule-based question formulation method. For each of the question type, we have the meta information from the local specialty annotation (Section 3.3). Then we design three to four templates for each of the question type. For example, for questions that ask about cuisine type, our templates include {CJK}UTF8gbsn
@@ Appendix B Question type and answer distribution
In Table 6, 7, and 8, we show concrete statistics about distribution of question types in each task. In Figure 10, we plot the distribution answer distribution for each of the question type in the tasks.
Figure 10: Answer distribution for each of the tasks.
Task Count
@@
Table 6: Distribution of text QA question types
Task Count
@@
Table 7: Distribution of single-image VQA question types
Task Count
@@
Table 8: Distribution of multi-image VQA question types
Appendix C Prompts used for evaluation
@@
Table 10: Prompts for zero-shot evaluation
Appendix D Interface of image collection, annotation and verification tool
@@
Appendix E More examples
E.1 Examples of the questions in the dataset

### v2

@@ 1 Introduction
Figure 2: The tasks in FoodieQA evaluate food culture understanding from three perspectives. Multi-image VQA requires the ability to compare multiple images, similar to how humans browse a restaurant menu. Single-image VQA assesses whether models can use visual information to better understand food culture. Text-based questions probe model performance without multimodal data. [^2] We only evaluate TextQA in Chinese to prevent bias introduced through translating dish names. The English translation is only for illustration purpose.Fine-grained attributes that the questions focus on are highlighted.
To build a regionally diverse dataset, we gather dishes and images selected by native Chinese speakers from various regions, covering 14 distinct cuisine types across China. To ensure the images used for benchmarking are fresh and have no chance of leaking into the pretraining data of VLMs, we collect images uploaded by local people, which are not publicly available online. We then define multiple attributes associated with the dishes and have native Chinese annotators create multiple-choice questions based on their expertise. Our dataset includes both vision-based question answering and text-based question answering tasks, as illustrated in Figure 2.
@@ 2 Related Work
In recent years, most food datasets have been designed for food image classification (Chen et al., 2017), food captioning (Ma et al., 2023), and recipe-focused generation and question answering (Yagcioglu et al., 2018; Min et al., 2018; Liu et al., 2022). For culture knowledge probing in the food domain, some of the recent datasets span multiple countries and include broad cultural or regional metadata (Min et al., 2018; Ma et al., 2023; Romero et al., 2024). However, they often use country as a proxy for culture, such as the country of origin for the food. For example, Palta and Rudinger (2023) introduced a test set to probe culinary cultural biases by considering US and non-US traditions, Zhou et al. (2024) construct a multicultural, multilingual dataset focusing on culinary knowledge, and Cao et al. (2024a) focuses on recipe transfer between Chinese and English. Investigating cultural differences within a country remains an under-explored area (Palta and Rudinger, 2023).
@@
Bugliarello et al. (2023) quantified the fine-grained vision-language understanding capabilities in existing models, focusing on aspects within the general domain. Later works focus on the culture understanding in VLMs Liu et al. (2023); Cao et al. (2024b). However, current fine-grained VL datasets (Zhang et al., 2021; Parcalabescu et al., 2022; Thrush et al., 2022; Hendricks and Nematzadeh, 2021) are often framed as binary classification tasks, which limits their difficulty. Concurrently with our work, Romero et al. (2024) and Nayak et al. (2024) have created culturally-diverse question-answering datasets across multiple countries. Our multi-choice vision question answering dataset that focuses on Chinese regional differences aims to advance the boundaries of fine-grained understanding in the context of food and culture.
@@ 3.3 Local Specialty Annotation
We also gather text annotations of representative local specialties for each cuisine type on our list. Annotators are asked to collect meta information for representative local dishes for each cuisine type, based on their life experience and knowledge obtained from the web. These meta-fields provide information beyond recipes, offering insights into how the food looks and tastes when people are eating it. An example is provided in Figure 4.
The 17 meta-info fields cover the appearance, taste, and culinary attributes of a dish. They include the food category, dish name, alternative names, main ingredient, characteristics of the main ingredient, three other key ingredients, dish flavor, presentation style, dish color, serving temperature (cold or warm), dishware used, region and province of origin, cuisine type, three primary cooking techniques, eating habits (if any), and reference links.
The annotation is done by eight native Chinese speakers, including five PhD students and three postdoctoral researchers from various provinces in China. [^6] The annotators are from Sichuan, Shaanxi, Guangdong, Jiangsu, Jiangxi, Shandong, and Chongqing. During the annotation process, we ensure that all collected data is either annotated or verified by individuals familiar with the local context. Specifically, annotators are assigned as follows: 1) They are asked to annotate local specialties for the cuisine types from their hometowns, guaranteeing that the annotations are provided by locals. 2) If a local annotator can not be found for a specific cuisine type, annotators are requested to seek assistance from friends who are from the respective region to verify or correct the metadata obtained from the web. Annotations in the following sections are conducted by the same annotators, if not mentioned otherwise.
@@ 3.4.1 Multi-image VQA
Once the questions and answers for the multi-image multiple-choice questions are collected, we verify the questions by asking the annotators (who did not create the questions) to answer them. If a question does not meet our defined criteria, annotators are instructed to flag it as a "bad question". Through this process, 87 questions were discarded. Additionally, when answering the questions, annotators are required to provide the rationale they use to reach the answer, as well as judge whether the question requires multi-hop reasoning. The user interface that we use for verification is shown in Figure 13. Each question is verified by two annotators, and we exclude the questions that do not have full agreement.
@@ 3.4.2 Single-Image VQA
Besides using images as multiple-choice answer options, we also ask diverse fine-grained questions about various aspects of a dish based on its meta-information (collected in Section 3.3). We identify dishes that have both meta-information annotations and collected images, and then create questions based on the meta-information. As shown in the example in Figure 2, the dish name is intentionally omitted from the questions to ensure they can only be answered by examining the visual features.
@@
We adopt a template-based approach, where a question about the same meta-field is asked multiple times, varying factors like the image of the dish, while the answer options are carefully selected from the wrong candidates in the meta-field to ensure that only one answer is correct. The single-image VQA questions are generated using a rule-based method, followed by thorough human verification that is similar to the multi-image VQA verification process. Please see details in the Appendix A.
@@
Similar to verification for the multi-image VQA questions, annotators are asked to answer the question given the text query and the corresponding image, and raise a "bad question" flag to filter out questions that does not satisfy the criteria. 88 questions were discarded as bad. Note that the name of the dish is not revealed in the text question so that the question needs to be answered based on visual information. Annotators are asked to write "I don’t know" in the rationale and randomly guess an answer if they think the question is beyond their knowledge.
@@ 3.5 Text Question Answering Annotation
We formulate the text-based questions by combining human annotations and rule-based generation. Similar to the single-image VQA approach described in Section 3.4.2, we generated questions and multiple-choice answer options based on the meta-information fields. However, instead of using the dish image, we included the dish name directly in the question. The questions are formulated using templates, where only the dish names and meta-fields are varied. A same human verification process to single-image question answering is included. 135 bad questions were discarded. Notice that annotators were asked to answer the questions based on their knowledge without using search engines, this makes the task challenging as it would be hard for one to answer questions about unfamiliar foods and regions without any other available information besides names of the food.
@@ 4.1 Human Validation
In Table 1, we calculate human accuracy and inter-annotator agreement scores based on human-verified questions, excluding those identified as bad questions. For the single-image VQA and text QA questions, given the diverse cultural backgrounds of the human annotators, some questions can be challenging if the required food culture knowledge falls outside an annotator’s cultural experience. For those questions, annotators are instructed to indicate "I don’t know" and randomly guess an answer, as one might not be familiar with all of the specific dishes or the fourteen cuisine types. These questions are marked as out-of-domain. Considering the randomly selected answers for these out-of-domain questions allow us to obtain lower bound agreement and human accuracy scores. [^7] Note that this is the only impact of the randomization. The ground truth label is annotated at an earlier stage of question formulation where the questions and choices are generated using the rule-based method. We also report Cohen’s Kappa ($\kappa$) and human accuracy separately for in-domain questions.
The human validation process involves three postdoctoral researchers and five PhD students who are native Chinese speakers as introduced in Section 3.3. Each question is verified and answered by two annotators who were not involved in the question formulation. We retain the out-of-domain questions for calculating human accuracy and later in evaluating model performance, as the lower agreement scores are only due to differences in the annotators’ cultural knowledge Plank (2022).
@@
Table 1: Statistics per task in FoodieQA.
@@
Figure 5: Region distribution of collected food images.
@@ 4.2 Image and Question Distribution
After human verification, we obtain 403 multi-image VQA questions, where each question needs to be answered with a set of four provided images. Single-image VQA tasks consists of 256 question in total, and text QA consists of 705 questions in total (Table 1). A considerable number of the VQA questions require multi-hop reasoning to predict the correct answer. We report the key statistics of the questions in Table 2. Please see more details in Appendix B.
@@ 5 Baselines: How Much of a Foodie are the LLMs/VLMs?
We evaluate open-weight and API-based state-of-the-art LLMs and VLMs to probe their culture knowledge in the food domain. We evaluate the models in both Chinese and English for the VQA tasks. The questions are translated to English using the DeepL free API [^8] https://www.deepl.com/en/translator and validated by two PhD students who are Chinese native speakers and fluent in English. To avoid bias in translating dish names, we conduct the TextQA task solely in Chinese.
@@ 5.1 Multi-Image VQA is Difficult
We experimented with four different prompts that utilized lists of images and texts or interleaved image-text inputs. Details can be found in Appendix D. As shown in Figure 6, when compared to the human accuracy of 91.69% in Chinese, the best-performing open-weight model, Idefics2-8B, achieves an accuracy of 50.87%, which is still significantly lower than human performance. This indicates that current state-of-the-art models are still weak at distinguishing differences among food from visual input. This underscores that multi-image understanding, especially in contexts requiring cultural knowledge in the food domain, remains a challenging problem. When evaluating on the translated English questions, model performance decreases for all models except Phi-3-vision.
@@ 5.2 Single-Image VQA Results
Table 3: Comparison of Multi-image and Single-image VQA Performance in Chinese and English. We report the best accuracy from four prompts. : see footnote.  [^10] Results with denote an estimate, calculated over 100 random samples, of human performance on the English Multi-Image and Single-Image VQA from one native speaker with no specialized knowledge of Chinese food culture.
@@ 5.3 Models are Strong at Text QA
From Figure 7, we see that the Qwen2-7B-instruct model surpasses human performance on the text QA task, where the questions are formulated based on the local specialty annotations in Section 3. Since the local specialty annotations are collected and summarized by local representatives, potentially incorporating information from public web resources such as Baidu-Baike, the high performance may be attributed to the inclusion of domain-specific training data.
@@ 6 Analysis
In this section, we explore which factors are important for fine-grained understanding of Chinese food culture.
Non-public images are crucial for fair evaluation.
We incorporate user-uploaded non-public images into our dataset to prevent data contamination during evaluation. To verify the importance of preserving these non-public images for fair evaluation, we compare model performance using web-sourced images instead. Specifically, we manually searched with dish names to obtain web images for 171 out of 256 questions in the Single-image VQA task. As shown in Table 4, replacing non-public images with web-sourced dish images made the task easier for baseline models, indicating potential data contamination from web sources. Therefore, the use of non-public images is crucial for ensuring fair evaluation.
Model non-public images web images
Qwen-VL-12B 43.75 47.95
Idefics2-8B 45.60 47.07
Yi-VL-6B 47.56 50.88
Table 4: Models obtain higher accuracy when evaluating with web images, which indicates possible data contamination. The accuracy scores are averaged over four prompts.
@@
In Single-image VQA, the default setting is to query with only dish image without specifying the dish name. We now examine whether the visual information is beneficial using the Idefics2-8B model. [^11] We selected this model because it supports text-only inputs, unlike some other models such as the Yi-VL series. Results are shown in Table 5, where we investigate two variants: querying the model with only the text question but revealing the dish name, versus providing both the dish image and the dish name. We observe that the Idefics2 model consistently performs better when dish images are available as visual clues. Please see comparison examples in Appendix F.2.
@@
(a) Accuracy by food categories.
(b) Accuracy by cuisine types.
@@ 7 Conclusion
Improved models or methods for understanding food culture may be essential for future progress in the FoodieQA challenge. Looking ahead, we aim to expand the dataset to include dishes from other countries and regions. Following Jacovi et al. (2023), we make our dataset a public benchmark on Huggingface at lyan62/FoodieQA with the CC BY-NC-ND 4.0 License. All of our data annotation and verification tools are freely available for re-use at github.com/lyan62/FoodieQA. We encourage the community to create Foodie datasets for their own language and culture groups.
@@ 8 Limitations
Acknowledgements
We are grateful to the volunteers for their generous contributions and efforts in providing high-quality food images that support our research. We extend our gratitude to Xi Liu, Yihe Zhang, Yu Sun, Yueyin Xu, Gefan Yang, Shixiong Wang, Penglong Ma, Daiwei Wang, Bo Cui, Yu Dong, Jinming Hu, Yufei Lin, and Zhongsheng Huang for serving as local experts. Their efforts in verifying and correcting the local specialty annotations and providing valuable feedback have been essential in ensuring the annotation’s accuracy and completeness. We also thank Fengyuan Liu, Ruixiang Cui, Zhi Zhang, Yu Sun, and many of our friends and family who helped spread the image collection survey on social media for wide regional and group coverage. Special thanks to Jordan Boyd-Graber and Jimmy Lin for providing helpful research advice. Wenyan Li is supported by the Lundbeck Foundation (BrainDrugs grant: R279-2018-1145) and a research grant (VIL53122) from VILLUM FONDEN. Jiaang Li is supported by Carlsberg Research Foundation (grant: CF221432). Li Zhou is supported by Shenzhen Science and Technology Research Fund (JCYJ20220818103001002) and Shenzhen Science and Technology Program (ZDSYS20230626091302006).
@@ Appendix A Rule-based question formulation
For text-based question answering we develop a rule-based question formulation method. For each question type, we have the meta information from the local specialty annotation (Section 3.3). Then we design three to four templates for each of the question type. For example, for questions that ask about cuisine type, our templates include {CJK}UTF8gbsn
@@ Appendix B Question type and answer distribution
In Table 78, and 9, we show concrete statistics about distribution of question types in each task. Figure 10 illustrates the answer distribution for questions categorized by type. Each horizontal bar independently displays the distribution of the answers regarding to the specific question type.
Figure 10: Answer distribution for each of the tasks. The questions are categorized by question type. Each color corresponds to a distinct answer, and each horizontal bar displays the distribution of these answers.
Question type Count
@@
Table 7: Distribution of text QA question types.
Question type Count
@@
Table 8: Distribution of single-image VQA question types .
Question type Count
@@
Table 9: Distribution of multi-image VQA question types .
Appendix C Annotation Cost and Compensation
Task Avg time/annotation Avg time/person
Local specialty collection 11.4 min/dish 10.3 hrs/person
Multi-image VQA question formulation 3.5 min/question 8.0 hrs/person
Multi-image VQA question verification 2.5 min/question 6.7 hrs/person
Single-image VQA verification 3.3 min/question 6.3 hrs/person
TextQA verification 1.2 min/question 5.7 hrs/person
Table 10: Average time per annotation and per person for annotation tasks.
In this work, the annotators are our colleagues who share co-authorship of the paper. This applies to the human annotation and validation process in Section 3.3, Section 3.4, and Section 4.1. The collection of images from private individuals, described in Section 3.2, was entirely voluntary and by community effort through the social platforms, WeChat and Douban.
The image collection period takes around one and a half months through the survey. Table 10 displays an estimation of the annotation time reported by annotators.
Appendix D Prompts used for evaluation
@@
Appendix E Interface of image collection, annotation and verification tool
@@
Appendix F More examples
F.1 Examples of the questions in the dataset